ANALYSIS AND FIT OF THE HIGH-RESOLUTION VISIBLE SPECTRUM OF THE TWO-EQUIVALENT-TOP MOLECULE BIACETYL

NOBUKIMI OHASHI, Department of Physics, Kanazawa University, Kakuma, Kanazawa 920-1192, Japan; \underline{JON T. HOUGEN}, Optical Technology Division, NIST, Gaithersburg, MD 20899-8441, USA; CHENG-LIANG HUANG, Department of Applied Chemistry, National Chiayi University, Chiayi, Taiwan; CHEN-LIN LIU, CHI-KUNG NI, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 106, P.O. Box 23-166, Taiwan.

The A^{1}A_{u}(S_{1}) − X^{1}A_{g}(S_{0}) LIF spectrum of biacetyl (CH_{3}C(=O)C(=O)CH_{3}) shows a long progression in the torsional vibrations of the two methyl tops. The S_{1} torsional level pattern can be qualitatively understood^{a} using local mode ideas applied to the two equivalent methyl rotors and symmetry ideas from the PI group G_{36}. For the present rotational analysis, we assigned a G_{36} symmetry species, two local-mode torsional quantum numbers, and rotational quantum numbers J_{Ka,Kc} to each observed torsion-rotation level. Transitions were globally fit with a two-equivalent-top computer program, which uses the molecular principal axis system, a free-rotor basis set for each top, a symmetric-top basis set for the rotational functions, and a single-step diagonalization procedure. The program is similar to our two-inequivalent-top program for N-methylacetamide.^{b} Last year, we refit 179 previously published^{c} rotational lines involving 7 torsional tunneling sublevels with zero or one quantum of torsional excitation in the S_{1} state of biacetyl, using 14 parameters (5 for S_{0}, 9 for S_{1}) to obtain a standard deviation of 0.0041 cm^{−1}. We have now extended that fit by adding 131 unpublished lines involving 5 tunneling sublevels with two quanta of torsional excitation, using 19 parameters to obtain a standard deviation of 0.0040 cm^{−1}. Including transitions with still higher torsional excitation leads to a degraded standard deviation. We thus suspect that a number of higher-order terms describing top-top-overall-rotation interactions are missing from the present model, but the possibility of misassignments in the higher-energy regions must also be considered.

aC.-L. Huang, C.-L. Liu, C.-K. Ni, and J. T. Hougen, J. Mol. Spectrosc. 233, 122-132 (2005).

bN. Ohashi, J. T. Hougen, R. D. Suenram, F. J. Lovas, Y. Kawashima, M. Fujitake, and J. Pyka, J. Mol. Spectrosc. 227, 28-42 (2004).

cC.-L. Huang, H.-H. Liu, C.-L. Liu, A. H. Kung, and C.-K. Ni, J. Chem. Phys. 117, 5165-5173 (2002).